# High-Throughput and Scalable Secure Inference Protocols for Deep Learning with Packed Secret Sharing

### v1

@@ Abstract
Most existing secure neural network inference protocols based on secure multi-party computation (MPC) typically support at most four participants, demonstrating severely limited scalability. Liu et al. (USENIX Security’24) presented the first relatively practical approach by utilizing Shamir secret sharing with Mersenne prime fields. However, when processing deeper neural networks such as VGG16, their protocols incur substantial communication overhead, resulting in particularly significant latency in wide-area network (WAN) environments. In this paper, we propose a high-throughput and scalable MPC protocol for neural network inference against semi-honest adversaries in the honest-majority setting. The core of our approach lies in leveraging packed Shamir secret sharing (PSS) to enable parallel computation and reduce communication complexity. The main contributions are three-fold: i) We present a communication-efficient protocol for vector–matrix multiplication, based on our newly defined notion of vector–matrix multiplication-friendly random share tuples. ii) We design the filter packing approach that enables parallel convolution. iii) We further extend all non-linear protocols based on Shamir secret sharing to the PSS-based protocols for achieving parallel non-linear operations. Extensive experiments across various datasets and neural networks demonstrate the superiority of our approach in WAN. Compared to Liu et al. (USENIX Security’24), our scheme reduces the communication upto $5.85\times$, $11.17\times$, and $6.83\times$ in offline, online and total communication overhead, respectively. In addition, our scheme is upto $1.59\times$, $2.61\times$, and $1.75\times$ faster in offline, online and total running time, respectively.
@@ I Introduction
Deep learning (DL) has emerged as a transformative methodology with widespread applications across diverse domains, particularly in computer vision and natural language processing. An increasing number of enterprises are deploying pre-trained DL models on computational servers to provide clients with inference services. However, this computational paradigm introduces significant privacy risks that warrant careful consideration. For example, the input data transmitted by clients may contain sensitive information, and DL models are often regarded as valuable assets that enterprises are reluctant to disclose. To address these dual privacy challenges, secure multi-party computation (MPC) is used to safeguard both input data and model privacy.
@@
MPC enables a set of $n$ parties to mutually compute some function on their private inputs without compromising the privacy of these inputs or the correctness of the outputs [13, 14, 15, 16]. In recent years, numerous studies have been conducted to enhance the communication and computational efficiency of MPC-based secure neural network inference. In the dishonest-majority setting, most existing protocols focus on the two-party scenario [17, 20, 21, 6, 18, 22, 23, 24, 25, 26]. In the honest-majority setting, most existing protocols consider the three-party scenario [9, 12, 27, 28, 29, 30] and four-party scenario [31, 32, 33, 34], where the protocols tolerate one corrupted party. Unfortunately, all the aforementioned protocols universally suffer from scalability limitations. This critical shortcoming was recently addressed by Liu et al. [5], who designed a scalable protocol that can support a large number of parties for secure neural network inference based on Shamir secret sharing [2] against semi-honest adversaries in the honest-majority setting. Nevertheless, when we rerun the scalable protocol on relatively deeper neural networks such as VGG16, it exhibits poor performance owing to significant communication overhead. Hence, It is necessary to develop a more communication-efficient protocol that can better accommodate the requirements of real-world scenarios.
In theory, Franklin and Yung introduced a parallelizable technique called packed Shamir secret sharing (PSS) [1] where $k$ secrets are packed into a single secret sharing. Its follow-up works [8, 3, 48] extend functionality for secure computation. Particularly, a new technique called sharing transformation is proposed in work [3]. The technique allows us to perform arbitrary linear maps on the secrets of PSS with optimal communication complexity. Based on these techniques, we are able to achieve parallel secure inference.
@@
Case I: For batch inference over $k$ input instances, naively upgrading the Shamir secret sharing scheme used in the protocol proposed by [5] to PSS could result in an amortized reduction in communication overhead by a factor of approximately $k$. However, in certain scenarios, such as a photo-based recognition application, only a single input may be available. If the above packed approach across multiple instances is still employed, it needs to introduce dummy inputs and pack them together with the real input to complete the inference process. More critically, this approach degenerates into a secure inference protocol similar to [5], thereby completely forfeiting the advantage of parallel computing provided by PSS.
@@
Case II: Given that the method described in Case I isn’t efficient for processing the single input, we can turn to employ the substantial parallelism in a single inference. To be specific, deep neural networks typically consist of linear layers and non-linear layers. Linear layers primarily execute matrix operations. For example, the fully connected layer is a vector-matrix multiplication. The convolutional layer can be expressed as a matrix multiplication [12], which can be achieved through multiple vector-matrix multiplications. All these operations can be efficiently implemented using blocked matrix algorithms, which are particularly amenable to parallel computing acceleration. Under the block-processing strategy, each block of the vector and matrix is encoded into an individual PSS value, and block-wise multiply-accumulate operations are performed using DN protocol based on PSS (a natural extension of the DN protocol in [4]). Fig. 1 shows a toy example.
Regarding non-linear layers, such as the ReLU function defined as $\text{ReLU}(x)=\text{Max}(0,x)$, the operations are element-wise. This means each output value depends solely on the single value at the corresponding position in the input tensor. Hence, the parallel non-linear operations are entirely achievable by packing $k$ values into an individual PSS value. With input and output always kept as PSS, consecutive layers are composable; consequently, the complete secure-inference pipeline is executed in parallel.
@@
I-A Practical Problem & Technique Overview
Nevertheless, a simple exploitation of parallelism in the inference process, as in Case II, is not practical, especially for linear layers.
@@ I-A Practical Problem & Technique Overview
First, the naive vector-matrix multiplication in Case II eliminates the benefit of using the PSS. As depicted in Fig. 1, each block of the vector is encoded into an individual PSS value, and block-wise multiply-accumulate operations are performed using DN protocol. Since the computed result remains in a partitioned form, this necessitates a sharing transformation to combine these partial results into the final output. Compared with Shamir secret sharing-based approach [5], the entire process instead incurs more communication overhead. The reason is that [5] only involves the DN protocol for vector-matrix multiplication.
@@
Second, when processing convolutional layers, zero padding of convolutional inputs incurs significant communication cost. As depicted in Case II, a convolutional layer can be achieved through multiple vector-matrix multiplications. In addition, zero padding is aslo frequently applied to the convolutional inputs in many neural networks, such as VGG16. Intuitively, we can directly achieve convolution computation by the naive vector-matrix multiplication. Unfortunately, when zero padding needs to be applied to convolutional inputs, it becomes necessary to combine all inputs with these zero padding values into new PSS values. This combination process incurs significant communication overhead.
@@
To address the first issue, inspired by the sharing transformation technique, we first define vector-matrix multiplication-friendly random share tuples (VM-RandTuple), which can be generated in batches by employing the Vandermonde matrix in the offline phase. Furthermore, in the DN protocol, the parties locally compute and send the result to $P_{1}$. Then, $P_{1}$ reconstructs the result and sends it back to the parties. A simple optimization to the second step is having $P_{1}$ combine these partial results into the final output and reshare the final output to the parties. Theoretically, integrating VM-RandTuple and the optimized DN protocol has the potential to achieve better online communication complexity vector-matrix multiplication.
@@
To address the second issue, we observed that the convolutional input is required to interact with every filter. Hence, the convolution process inherently involves parallel computation. In light of this, we propose a filter packing approach, i.e., we pack $k$ values that are selected respectively from the same position of $k$ filters into a single packed secret share. Correspondingly, we also pack $k$ copies of the input into a single packed secret share. Then based on the packing approach, we achieve completing the convolution computation in parallel. Further, when zero padding is applied to the convolutional input without communication. Finally, since $k$ copies of the input are packed into a single packed secret share, we also carefully account for the inter-layer connectivity by designing some protocols.
@@ I-B Contributions
Building upon the above techniques, in this paper, we propose a parallel and scalable MPC protocol by leveraging PSS for secure inference, where the model from companies and input data from clients are secret-shared among all computational servers. The main contributions are summarized below.
  • Efficient Vector-matrix Multiplication. We define vector-matrix multiplication-friendly random share tuples and design a protocol to generate these tuples for computing vector-matrix multiplication. We further combine these tuples with the DN [4] protocol, resulting in an efficient protocol for vector-matrix multiplication with a single round of online communication.
@@
  • Parallel Non-linear Operations. We extend all protocols based on Shamir secret sharing for non-linear operations in [5] to the PSS-based protocols, which enable parallel computation of the non-linear operations, resulting in efficient protocols for parallel ReLU with $O(\text{log}_{2}{\ell})$ rounds of online communication.
  • Efficient Secure Inference. We implement all protocols and conduct extensive secure inference experiments on various neural networks in both LAN and WAN settings with varying numbers of parties and different packaging parameters. Experimental results demonstrate progressively enhanced performance advantages with increasing numbers of participating entities and deeper neural network architectures. Compared to Liu et al. [5], our scheme reduces the communication upto $5.85\times$, $11.17\times$, and $6.83\times$ in offline, online, and total communication overhead, respectively. At the same time, our scheme is upto $1.59\times$, $2.61\times$, and $1.75\times$ in offline, online and total running time, respectively. Furthermore, experimental results reveal the superiority of our protocol in terms of memory overhead. For example, we simulated 63 participants on 11 computational servers, each of which was equipped with 64GB of RAM. Our protocol reliably supports VGG16 inference. In contrast,  [5] suffers from memory overflow problems.
@@ II-A Notation
Let $p$ be a Mersenne prime, which is the form of $p=2^{\ell}-1$ for some prime $\ell$. In secure neural network inference, $\ell$ is chosen to be 31 and 61, which is sufficient. $\mathbb{F}_{p}$ denotes a finite field. We denote by $[n]$ the set $\{0,1,\cdots,n-1\}$, where $n\in\mathbb{N},\mathbb{N}$ is the set of natural numbers. We use the lowercase letter $x$ to denote a scalar and $x_{i}$ to denote the $i$-th bit of the scalar $x$. The bold lowercase letter $\boldsymbol{x}=\{\boldsymbol{x_{i}}\}_{i=0}^{n-1}$ is a vector, $\boldsymbol{x}_{i}$ is the $i$-th element of the vector $\boldsymbol{x}$, $\boldsymbol{x}_{i,j}$ is $j$-th bit of the $i$-th element of the vector $\boldsymbol{x}$ and $\boldsymbol{x}^{i}$ is indexed by $i$ to distinguish different vectors. Correspondingly, uppercase letter $X$ is a matrix and $X_{i,j}$ is the element indexed by row $i$ and column $j$ of the matrix $X$. The bold uppercase letter $\boldsymbol{X}$ is a tensor. Van(n,nt)Van(n,n-t) is an $n$-by-(nt)(n-t) Vandermonde matrix.
@@ II-B Packed Shamir Secret Sharing Scheme
  • Multiplicative: For all $d_{1},d_{2}\geq k-1,d_{1}+d_{2}<n$ and for all $\boldsymbol{x},\boldsymbol{y}\in\mathbb{F}_{q}^{k},\llbracket\boldsymbol{x}\rrbracket_{d_{1}}\times\llbracket\boldsymbol{y}\rrbracket_{d_{2}}=\llbracket\boldsymbol{x}*\boldsymbol{y}\rrbracket_{d_{1}+d_{2}}$. If $d_{1}=d_{2}=\frac{n-1}{2}$, we can compute $\llbracket\boldsymbol{x}\rrbracket_{d}*\llbracket\boldsymbol{y}\rrbracket_{d}=\llbracket\boldsymbol{x}\times\boldsymbol{y}\rrbracket_{d}$ by a natural extension of the DN multiplication protocol in [4]. We summarize the functionality $\mathcal{F}_{Pmult-DN}$ in Appendix -A. The protocol is as follows:
@@ II-C Neural Network
Non-Linear Layer (ReLU and Pool). In deep learning, the non-linear layers consist of activation functions and pooling functions. The activation functions are applied in each element of input independently. The common activation functions is ReLU function: $\text{ReLU}(x)=\text{Max}(0,x)$. The pooling function reduces the output size. One of the most popular pooling functions in CNN is the MaxPool function that arranges inputs into several windows and calculates the maximum for each window. We construct the parallel non-linear operators in §V.
@@ II-D Model
We use an outsourcing computing model for secure inference. We assume there are a client, a model owner and $n\in\mathbb{N}$ computational servers, denoted by $P_{1},P_{2},\cdots,P_{n},$ where $n=2d+1,d\in\mathbb{N}$. When a client wants to perform secure inference, the client and model owner respectively distribute their own data across $n$ servers by a degree-$d$ PSS scheme. Then $n$ servers complete a secure inference and send the result shares to the client. Finally, the client can get the inference result by reconstructing the shares.
We describe our protocols in the semi-honest model with passive security. The client and model owner are honest. The number of corrupted servers that are assumed to be poly-time bounded is at most t=dk+1t=d-k+1 for a natural extension of the DN [4] multiplication protocol, where $2\leq k\leq d$. If $k=1$, the PSS becomes standard Shamir secret sharing. The client and the model owner do not collude with the servers. It is worth noting that a degree-$d$ PSS can pack $k\in\mathbb{N}$ secrets, and $k$ and default positions $\{\boldsymbol{s}_{i}\}_{i=0}^{k-1}$ are publicly known to the client, model owner and servers.

### v2

# High-Throughput and Scalable Secure Computation Protocols for Deep Learning with Packed Secret Sharing

@@ Abstract
Most existing privacy-preserving deep learning protocols based on secure multi-party computation (MPC) typically support at most four participants, demonstrating severely limited scalability. Liu et al. (USENIX Security’24) presented the first relatively practical approach by utilizing Shamir secret sharing with Mersenne prime fields. However, when processing deeper neural networks such as VGG16, their protocols incur substantial communication overhead, resulting in particularly significant latency in wide-area network (WAN) settings. To tackle this challenge, in this paper, we propose a high-throughput and scalable MPC protocol for deep learning in the honest-majority setting. The core of our approach lies in leveraging packed Shamir secret sharing (PSS) to enable parallel computation and reduce communication complexity. The main contributions are three-fold: i) We define a primitive called vector-matrix multiplication-friendly random share tuples and present a communication-efficient protocol for vector–matrix multiplication based on the primitive. ii) Inspired by the inherent parallelizability of multiple-input multiple-output convolutions, we design the filter packing approach that enables parallel convolution. iii) We further extend all non-linear protocols based on Shamir Secret Sharing to their PSS variants, which enable parallel non-linear evaluations. Extensive experiments across various datasets and neural networks demonstrate the superiority of our approach in WAN. Specifically, compared to Liu et al. (USENIX Security’24), our scheme reduces the communication by up to $5.9\times$, $11.2\times$, and $6.8\times$ in offline, online and total communication overhead, respectively. In addition, our scheme is up to $1.6\times$, $2.6\times$, and $1.8\times$ faster in offline, online and total running time, respectively.
@@ I Introduction
Deep learning (DL) has emerged as a transformative methodology with widespread applications across diverse domains, particularly in computer vision and natural language processing. However, its reliance on large-scale data, which often includes sensitive information, raises significant privacy concerns that require careful consideration. Secure multi-party computation (MPC) offers a promising solution to this challenge by enabling a set of $n$ parties to jointly compute a function over their private inputs while preserving both input privacy and output correctness [1, 2, 3, 4]. Integrating MPC with deep learning thus allows multiple parties to collaboratively perform training or inference tasks without exposing their private data.
@@
In recent years, numerous studies have been conducted to enhance the communication and computational efficiency of MPC-based privacy-preserving deep learning (PPDL) protocols. In the dishonest-majority setting, most existing protocols focus on the two-party scenario [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. In the honest-majority setting, most existing protocols consider the three-party scenario [15, 16, 17, 18, 19, 20] and four-party scenario [21, 22, 23, 24], where the protocols tolerate one corrupted party. Unfortunately, all the aforementioned protocols universally suffer from scalability limitations. This critical shortcoming was recently addressed by Liu et al. [25], who designed a scalable protocol that can support a large number of parties for PPDL based on Shamir secret sharing [26] against semi-honest adversaries in the honest-majority setting. Nevertheless, when we rerun the scalable protocol on relatively deeper neural networks such as VGG16, it exhibits poor performance owing to significant communication overhead. Hence, it is necessary to develop a more communication-efficient protocol that can better accommodate the requirements of real-world scenarios. For ease of reference, we denote this work [25] as “LXY24”.
@@
In theory, Franklin and Yung introduced a parallelizable technique called packed Shamir secret sharing (PSS) [27], which packs $k$ secrets into a single secret sharing. Naively replacing the Shamir secret sharing scheme in the LXY24 with a PSS scheme can enable parallel processing of $k$ instances, which results in an amortized reduction in communication overhead by a factor of approximately $k$. However, in certain scenarios, such as a photo-based recognition application, only a single input may be available. If the above packed approach across multiple instances is still employed, it needs to introduce dummy inputs and pack them together with the real input to complete the inference process. More critically, this approach degenerates into the protocol similar to LXY24, thereby completely forfeiting the advantage of parallel computing provided by PSS.
@@
Given that the naive method isn’t efficient for processing a single input, we can turn to employ the substantial parallelism within a single inference pass. To be specific, deep neural networks typically consist of linear layers and non-linear layers. Linear layers primarily execute matrix operations. For example, the fully connected layer is a vector-matrix multiplication. The convolutional layer can be expressed as a matrix multiplication [16], which can be achieved through multiple vector-matrix multiplications. All these operations can be efficiently implemented using the blocked matrix strategy, which is particularly amenable to parallel computing acceleration. Under the block-processing strategy, each block of the vector and matrix is encoded into an individual PSS value, and block-wise multiply-accumulate operations are performed using DN protocol based on PSS (a natural extension of the DN protocol in [28]). Fig. 1 shows a toy example. Particularly,  [29] proposes the sharing transformation protocol that can transform a PSS packing secrets $(x_{1},x_{2})$ into one packing $(x_{1}+x_{2},0)$ for the case $k=2$. Regarding non-linear layers, such as the ReLU function defined as $\text{ReLU}(x)=\text{Max}(0,x)$, the operations are element-wise. This means each output value depends solely on the single value at the corresponding position in the input tensor. Hence, the parallel non-linear operations are entirely achievable by packing $k$ values into an individual PSS value. Nevertheless, the above simple exploitation of parallelism in a single inference pass is not practical, especially for linear layers.
@@
First, the trivial vector-matrix multiplication eliminates the benefit of using the PSS. As depicted in Fig. 1, each block of the vector is encoded into an individual PSS value, and block-wise multiply-accumulate operations are performed using the DN protocol. Since the computed result remains in a partitioned form, this necessitates an interactive sharing transformation protocol to combine these partial results into the final output. Compared with LXY24, the entire process instead incurs more communication overhead. The reason is that LXY24 only involves the DN protocol for vector-matrix multiplication.
@@
Second, when processing convolutional layers, zero padding of convolutional inputs incurs significant communication cost. Recall that zero padding is also frequently applied to the convolutional inputs in many neural networks, such as VGG16. A convolutional layer can be achieved through multiple vector-matrix multiplications. Unfortunately, when zero padding needs to be applied to convolutional inputs, it becomes necessary to combine all inputs with these zero padding values into new PSS values. This combination process incurs significant communication overhead.
I-A Technique Overview
@@
To address the first issue, we define a primitive called vector-matrix multiplication-friendly random share tuples (VM-RandTuple), which can be generated in batches by employing the Vandermonde matrix in the offline phase. Recall that the DN protocol [28] consists of two steps: First, the parties locally compute and send the results to $P_{1}$. Second, $P_{1}$ reconstructs the result and sends it back to the parties. Our optimization is motivated by the fact that $P_{1}$ holds the partial results plaintext, and the final output arises as a composition of partial results and VM-RandTuple, enabling the interactive sharing transformation protocol to be delegated and locally implemented within $P_{1}$. Theoretically, our optimization has the potential to achieve better online communication complexity vector-matrix multiplication.
@@
To address the second issue, we observed that the input is convolved with each filter in either a single-input multiple-output or multiple-input multiple-output convolution process. Therefore, such a convolution process inherently involves parallel computation. In light of this, we propose a filter packing approach, i.e., we pack $k$ values that are selected respectively from the same position of $k$ filters into a single packed secret share. Correspondingly, we also pack $k$ copies of each input pixel into a single packed secret share. Based on this packing approach, not only do we achieve a parallel convolution computation, but also there is no need to recombine zero padding values with the input to form new PSS representations. Instead, we directly pack $k$ copies of zero padding values locally. Moreover, due to $k$ copies of the input being packed into a single packed secret share, which differs from the packing approach of the vector-matrix multiplication, we need to design a protocol for the inter-layer connectivity.
@@ I-B Contributions
Building upon the above techniques, in this paper, we propose a parallel and scalable MPC protocol by leveraging PSS for PPDL. The main contributions are summarized below.
  • Efficient Vector-matrix Multiplication. We define a primitive called vector-matrix multiplication-friendly random share tuples and design a protocol to generate these tuples for computing vector-matrix multiplication. We further combine these tuples with the DN [28] protocol, resulting in an efficient protocol for vector-matrix multiplication with a single round of online communication.
@@
  • Parallel Non-linear Operations. We extend all protocols based on Shamir secret sharing for non-linear operations in LXY24 to the PSS-based protocols, which enable parallel computation of the non-linear operations, resulting in efficient protocols for parallel ReLU with $O(\text{log}_{2}{\ell})$ rounds of online communication.
  • Evaluations. We implement all protocols and conduct extensive evaluations on various neural networks in both LAN and WAN settings with varying numbers of parties and different packaging parameters. Experimental results demonstrate progressively enhanced performance advantages with increasing numbers of participating entities and deeper neural network architectures. Compared to LXY24, our scheme reduces the communication by up to $5.9\times$, $11.2\times$, and $6.8\times$ in offline, online, and total communication overhead, respectively. At the same time, our scheme is up to $1.6\times$, $2.6\times$, and $1.8\times$ in offline, online and total running time, respectively. Furthermore, experimental results reveal the superiority of our protocol in terms of memory overhead. For example, in our experiment setup, our protocol reliably supports VGG16 inference, whereas LXY24 suffers from memory overflow problems.
@@ II-A Notation
Let $p$ be a Mersenne prime, which is the form of $p=2^{\ell}-1$ for some prime $\ell$. In PPDL, $\ell$ is chosen to be 31 and 61, which is sufficient. $\mathbb{F}_{p}$ denotes a finite field. We denote by $[n]$ the set $\{0,1,\cdots,n-1\}$, where $n\in\mathbb{N},\mathbb{N}$ is the set of natural numbers. We use the lowercase letter $x$ to denote a scalar and $x_{i}$ to denote the $i$-th bit of the scalar $x$. The bold lowercase letter $\boldsymbol{x}=\{\boldsymbol{x_{i}}\}_{i=0}^{n-1}$ is a vector, $\boldsymbol{x}_{i}$ is the $i$-th element of the vector $\boldsymbol{x}$, $\boldsymbol{x}_{i,j}$ is $j$-th bit of the $i$-th element of the vector $\boldsymbol{x}$ and $\boldsymbol{x}^{i}$ is indexed by $i$ to distinguish different vectors. Correspondingly, uppercase letter $X$ is a matrix and $X_{i,j}$ is the element indexed by row $i$ and column $j$ of the matrix $X$. The bold uppercase letter $\boldsymbol{X}$ is a tensor. Van(n,nt)Van(n,n-t) is an $n$-by-(nt)(n-t) Vandermonde matrix.
@@ II-B Packed Shamir Secret Sharing Scheme
  • Multiplicative: For all $d_{1},d_{2}\geq k-1,d_{1}+d_{2}<n$ and for all $\boldsymbol{x},\boldsymbol{y}\in\mathbb{F}_{q}^{k},\llbracket\boldsymbol{x}\rrbracket_{d_{1}}\times\llbracket\boldsymbol{y}\rrbracket_{d_{2}}=\llbracket\boldsymbol{x}*\boldsymbol{y}\rrbracket_{d_{1}+d_{2}}$. If $d_{1}=d_{2}=\frac{n-1}{2}$, we can compute $\llbracket\boldsymbol{x}\rrbracket_{d}*\llbracket\boldsymbol{y}\rrbracket_{d}=\llbracket\boldsymbol{x}\times\boldsymbol{y}\rrbracket_{d}$ by a natural extension of the DN multiplication protocol in [28]. We summarize the functionality $\mathcal{F}_{PDN}$ in Appendix -C. The protocol is as follows:
@@ II-C Neural Network
Non-Linear Layer (ReLU and Pool). In deep learning, the non-linear layers consist of activation functions and pooling functions. The activation functions are applied to each element of input independently. The common activation functions is ReLU function: $\text{ReLU}(x)=\text{Max}(0,x)$. The pooling function reduces the output size. One of the most popular pooling functions in CNN is the MaxPool function that arranges inputs into several windows and calculates the maximum for each window. We construct the parallel non-linear operators in §V.
@@ II-D Model
We assume there are $n\in\mathbb{N}$ parties, denoted by $P_{1},P_{2},\cdots,P_{n},$ where $n\geq 2d+1,d\in\mathbb{N}$. We describe our protocols in the semi-honest model with passive security. For every pair of parties, assume there exists a private and authentic synchronous channel. The number of corrupted parties that are assumed to be poly-time bounded is at most t=dk+1t=d-k+1 for a natural extension of the DN [28] multiplication protocol, where $2\leq k\leq d$. If $k=1$, the PSS becomes standard Shamir secret sharing. It is worth noting that a degree-$d$ PSS can pack $k\in\mathbb{N}$ secrets, and $k$ and default positions $\{\boldsymbol{s}_{i}\}_{i=0}^{k-1}$ are publicly known to the $n\in\mathbb{N}$ parties.